What is the probability of flipping heads and rolling an odd number?
These would be independent events; therefore, we can multiply the probabilities of each of the two events.
Probability of flipping a head: 1/2
Probability of rolling an odd number with a single die: 1/6
Required probability : 1/2 x 1/6 = 1/12
There are (10 x 9)/2 = 45 different possible pairs of 2 teachers. For each of these . . .
There are (30 x 29)/2 = 435 different possible pairs of students.
The total number of different committees that can be formed is (45 x 435) = 19,575 .
How do you see that the sampling distribution of the mean is normal?
According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.
The necessary requirements are shown in bold.
According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.
The necessary requirements are shown in bold.
According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.
The necessary requirements are shown in bold.
According to the Central Limit Theorem, the mean of a sufficiently large number of independent random variables which have a well defined mean and a well defined variance, is approximately normally distributed.
The necessary requirements are shown in bold.
When two dice are rolled find the probability of getting a sum less than 10?
There are 36 combinations of two dice. each have 1 to 6
To get 10 or more requires 4 +6, 5 +5, 5 +6 , 6+5, 6 +4, or 6+6, which are 6 combinations, leaving 36-6 = 30 combinations of less than 10
30/36 = 5/6 probability of getting sum less than 10
What is the probability that z is greater than .5?
Pr(Z > 0.5) = 0.3085 (approx).
Pr(Z > 0.5) = 0.3085 (approx).
Pr(Z > 0.5) = 0.3085 (approx).
Pr(Z > 0.5) = 0.3085 (approx).
What is the probability that a randomly selected person will be both over 50 and female?
The answer depends on the demography of the population from which the person is randomly selected.
The answer depends on the demography of the population from which the person is randomly selected.
The answer depends on the demography of the population from which the person is randomly selected.
The answer depends on the demography of the population from which the person is randomly selected.
What is the probability of getting 4 red cards from a standard deck of 52 cards?
If you are picking a card and putting it back before picking the next, it would be 1/2 * 1/2 * 1/2 * 1/2 = 1/16
If you are not replacing the cards, it would be 26/52 * 25/51 * 24/50 * 23/49 = 358800 / 6497400, or approximately 5.5%.
How many ways word crisis arranged?
There are 6 letter in the word crisis so the answer seems to be 6! (6 factorial), but there are two 'i's and two 's's which are indistinguishable and each of which can be arranged in 2! ways. So:
number_of_ways = 6!/2!2!
= 180 ways.
The probability of choosing a blue marble is 5 in 15 or 1 in 3. The probability of then choosing a green marble is 5 in 14. (One is missing) Multiply the two probabilities and you get 5 in 42.(P = 0.1190... about 12%).
What is the probability of rolling either a 5 or 6 or 8 or 9 before rolling a 7 with two dice?
Consider the three events:
A = rolling 5, 6, 8 or 9.
B = rolling 7
C = rolling any other number.
Let P be the probability of these events in one roll of a pair of dice.
Then P(A) = P(5) + P(6) + P(8) + P(9) = 18/36 = 1/2
P(B) = P(7) = 6/36 = 1/6
and P(C) = 1 - [P(A) + P(B)] = 1/3
Now P(A before B) = P(A or C followed by A before B)
= P(A) + P(C)*P(A before B)
= 1/2 + 1/3*P(A before B)
That is, P(A before B) = 1/2 + 1/3*P(A before B)
or 2/3*P(A before B) = 1/2
so that P(A before B) = 1/2*3/2 = 3/4
How many possible ways to arrange the word Mississippi?
Nonew, since mississipi is not a word!
Nonew, since mississipi is not a word!
Nonew, since mississipi is not a word!
Nonew, since mississipi is not a word!
What are the differences between probability sampling and non probability sampling?
In a probability sample, each unit has the same probability of being included in the sample. Equivalently, given a sample size, each sample of that size from the population has the same probability of being selected.
This is not true for non-probability sampling.
How do you find the mean of frequency distributions?
For discrete distributions, suppose the variable X takes the specific value x with probability P(X=x)
Then add together x * P(X = x) for all possible values of x.
For continuous distributions, suppose the probability distribution function of the variable X is f(x).
Then the mean is the integral of x*f(x) with respect to x, taken over all possible values of x.
How many equally likely outcomes are in a 6 faced dice?
Six of them, if the cube is honest (balanced).
What is the probability that a family with 4 children has at least 2 girls?
Assuming that boys and girls are equally likely, it is 11/16.
What is the probability of it raining in spring?
The answer depends on where on earth the question is about.
The answer depends on where on earth the question is about.
The answer depends on where on earth the question is about.
The answer depends on where on earth the question is about.
Pr(4 tails | at least 3 tails) = Pr(4 tails and at least 3 tails)/Pr(at least 3 tails)
= Pr(4 tails)/Pr(at least 3 tails)
= (5/32) / (1/2) = 5/16.
Pr(4 tails | at least 3 tails) = Pr(4 tails and at least 3 tails)/Pr(at least 3 tails)
= Pr(4 tails)/Pr(at least 3 tails)
= (5/32) / (1/2) = 5/16.
Pr(4 tails | at least 3 tails) = Pr(4 tails and at least 3 tails)/Pr(at least 3 tails)
= Pr(4 tails)/Pr(at least 3 tails)
= (5/32) / (1/2) = 5/16.
Pr(4 tails | at least 3 tails) = Pr(4 tails and at least 3 tails)/Pr(at least 3 tails)
= Pr(4 tails)/Pr(at least 3 tails)
= (5/32) / (1/2) = 5/16.
What is the probability of rolling a factor of 2 on a die?
There are three factors of 2 on a regular six-sided die. Therefore, it is 3/6, or 1/2.
What is the probability of an event b that impossible?
The probability of an impossible event is 0.
The probability of an impossible event is 0.
The probability of an impossible event is 0.
The probability of an impossible event is 0.
How many number 7 in a deck of 52 cards?
There are four 7s, which are the 7 of club, 7 of heart, 7 of spade and the 7 of diamond. Each number from 1 to 10 including the Jack, Queen and King have their cards of club, heart, spade and diamond each, which makes 13 (1 to 10 plus J Q and K) x 4 (club, heart, spade and diamond) = 52 cards
Can a negative z score be greater than a positive z score?
Assume the z-score is relative to zero score. In simple terms, assume that we have 0 < z < z0, where z0 is the arbitrary value.
Then, a negative z-score can be greater than a positive z-score (yes). How? Determine the probability of P(-2 < z < 0) and P(0 < z < 1). Then, by checking the z-value table, you should get:
P(-2 < z < 0) ≈ 0.47725
P(0 < z < 1) ≈ 0.341345
How does sample size affect the size of your standard error?
The standard error should decrease as the sample size increases. For larger samples, the standard error is inversely proportional to the square root of the sample size.
The standard error should decrease as the sample size increases. For larger samples, the standard error is inversely proportional to the square root of the sample size.
The standard error should decrease as the sample size increases. For larger samples, the standard error is inversely proportional to the square root of the sample size.
The standard error should decrease as the sample size increases. For larger samples, the standard error is inversely proportional to the square root of the sample size.
What situation represents the highest entropy of a deck of playing cards?
stacking the cards in order of face value.....(apex)
Is the uniform probability distribution is symmetric about the mode?
Yes, the uniform probability distribution is symmetric about the mode. Draw the sketch of the uniform probability distribution. If we say that the distribution is uniform, then we obtain the same constant for the continuous variable.
* * * * *
The uniform probability distribution is one in which the probability is the same throughout its domain, as stated above. By definition, then, there can be no value (or sub-domain) for which the probability is greater than elsewhere. In other words, a uniform probability distribution has no mode. The mode does not exist.
The distribution cannot, therefore, be symmetric about something that does not exist.